7. A 58 kg runner is traveling at a velocity of 5 m/s. What is her momentum (in Ns)?
Answer #1

Mass, m= 58kg

Velocity, v= 5 m/s

So, her momentum,

p= m*v = 58*5 = 290 Ns (Answer)